OpenStudy (mackenzie2013):

Find the principal amount invested if the amount in a semi-annually compounded account with a 5.05% interest rate is $16,360.93 after 60 months.

OpenStudy (anonymous):

A = 16,360.93(1+0.55/2)^4(60)

OpenStudy (anonymous):

\[a = 16360.93(1+\frac{ .55 }{ 2 })\]^240

OpenStudy (anonymous):

Not too good with typing these out with this format.